OpenStudy (anonymous):

Solve: -2(2x + 5) - 3 = -3(x - 1) 16 -16 -10 -13

HanAkoSolo (jamierox4ev3r):

just distribute everything. For example, on the first one.. -2(2x+5) = -4x=10 Can you do the other one, than solve? @MyChem ?

OpenStudy (anonymous):

?

HanAkoSolo (jamierox4ev3r):

distribute so that you can get rid of the parentheses. -2*2x=-4x and -2*5=-10. its just multiplication. Can you do the other for me @MyChem ?

HanAkoSolo (jamierox4ev3r):

I want to make sure you understand, lol

OpenStudy (anonymous):

lol alright the ( x-1) part you want me to do?

HanAkoSolo (jamierox4ev3r):

yes what does -3(x-1) equal after you distribute it?

OpenStudy (anonymous):

this ones too long can we do a different one? :p

HanAkoSolo (jamierox4ev3r):

well its really not that difficult once you solve, just simplify by distribution, then make sure that you solve for both sides. Would you be able to find the answer if I set up the equation for you? -4x-10=-3x+3 ^ now solve that equation right there. it is a simplified form

HanAkoSolo (jamierox4ev3r):

@MyChem come back here!! lol :P I will help you with the other one too ^_^